OpenStudy (anonymous):

Alexei wants to hang a mirror in his boat but the mirror and frame must not have an area larger than 19.25 square feet. The mirror is 3 feet wide and 5 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The border must have equal width on each side.)

OpenStudy (anonymous):

1. 4x2 + 16x - 4.25 = 0 2. 3x2 + 5x - 19.25 = 0 3 4x2 + 16x + 15 = 0 4 2x2 + 8x - 19.25 = 0

jimthompson5910 (jim_thompson5910):

Let x = width of the border The old length is 5 feet. The new length is 5+2x feet long. We're adding 2x to represent the idea that we're adding a border with a width of x feet to each side of the length The same idea applies to the width, the width goes from 3 feet to 3+2x feet. The total area is then A = LW A = (5+2x)(3+2x) A = 15 + 16x + 4x^2 A = 4x^2 + 16x + 15 Since this area cannot exceed 19.25 square feet, we know that the max area is 19.25 square feet So if the area is maxed out then 19.25 = 4x^2 + 16x + 15 From here, get everything to one side and simplify 19.25 = 4x^2 + 16x + 15 0 = 4x^2 + 16x + 15-19.25 0 = 4x^2 + 16x - 4.25 4x^2 + 16x - 4.25 = 0 So the answer is choice A